OpenStudy (anonymous):

What is the net ionic equation of the reaction of BeCl2 with NaOH? Express you answer as a chemical equation including phases.

thomaster (thomaster):

first write down the complete ionic equation, then get rid of the spectator ions and you'll have the net ionic equation :)

OpenStudy (paxpolaris):

\[BeCl_2\left( aq \right)+2NaOH \left( aq \right)\to2NaCl \left( aq \right)+Be \left( OH \right)_2(s)\]

thomaster (thomaster):

@PaxPolaris I don't remember asking you to first answer the question I asked to the topic starter, and second to give away the entire answer without any explanation or guidance...